# A Hidden Pulse: Uncovering a New Timing Signal in Cygnus X-1 with AstroSat

### v1

@@ Abstract
The study of fast variability properties in X-ray binaries advances our understanding of the physical processes and geometric properties of the accretion flow around the compact object. In this work, we study the evolution of the timing properties of Cygnus X-1 with AstroSat/LAXPC, during the transition of the source from the hard to soft state in 2017. We use a novel frequency-segmented technique to fit simultaneously the cross spectra and part of the power spectra and coherence function with a multi-Lorentzian model and predict the phase-lags and the complementary parts of the power spectra and coherence function. We identify previously undetected variability components, one of which manifests as a narrow dip in the coherence function and a broad drop in the phase-lag spectrum at the same frequency. This dip in coherence, which we detected for the first time in Cygnus X-1 at energies above 3 keV, appears in a state in which the source shows high-amplitude radio variability and significant hard X-ray polarization. While the contribution of the compact jet in X-rays is debated in the literature, this study provides a new avenue for investigating jet properties as well as the geometry of the Comptonizing medium.
@@ 4 Discussion
The two observations (1180 and 1210), where we detect the dip in coherence, lie in the HI state, concurrent with significant polarization measurements and high radio variability. Not only that, the four observations (1094, 1122, 1180, and 1210), which require the additional Lorentzians - beyond the five main components - also occur in the HI state. It is important to note here that the change in coherence pattern and their association with spectral states is not restricted to hidden QPOs (e.g., Fogantini et al., 2025). Alabarta et al. (2025) found that in MAXI J1348 there is an increase in radio emission during the epoch when the Type-C QPOs resulted in a dip in the coherence. Even the non-linear variability patterns at the QPO frequencies, explored through bicoherence studies, have been found to be correlated with spectral states and jet properties in BH LMXBs (see e.g., Arur & Maccarone, 2022, and references therein). These findings suggest that, apart from the main Comptonizing medium, the base of the jet could also act as an additional medium for Comptonization. Such hybrid geometries for the corona have been suggested by some recent spectral-timing studies (e.g., García et al., 2021; Méndez et al., 2022; Rout et al., 2023).

### v2

@@ Abstract
The study of fast variability properties in X-ray binaries advances our understanding of the physical processes and geometric properties of the accretion flow around the compact object. In this work, we study the evolution of the timing properties of Cygnus X-1 with AstroSat/LAXPC, during the transition of the source from the hard to soft state in 2017. We use a novel frequency-segmented technique to fit simultaneously the cross spectra and parts of the power spectra and coherence function with a multi-Lorentzian model and predict the phase-lags and the complementary parts of the power spectra and coherence function. We study the evolution of the frequency and power of the main variability components that are present throughout all the states. In particular, we identify previously undetected variability components, one of which manifests as a narrow dip in the coherence function and a broad drop in the phase-lag spectrum at the same frequency. This dip in coherence, which we detected for the first time in Cygnus X-1 at energies above 3 keV, appears in a state in which the source shows high-amplitude radio variability and significant hard X-ray polarization. While the contribution of the compact jet in X-rays is debated in the literature, this study provides a new avenue for investigating jet properties as well as the geometry of the Comptonizing medium.
@@ 4 Discussion
The two observations (1180 and 1210), where we detect the dip in coherence, lie in the HI state, concurrent with significant polarization measurements and high radio variability. Not only that, the four observations (1094, 1122, 1180, and 1210), which require the additional Lorentzians - beyond the five main components - also occur in the HI state. It is important to note here that the change in coherence pattern and their association with spectral states is not restricted to hidden QPOs (e.g., Fogantini et al., 2025). Alabarta et al. (2025) found that in MAXI J1348 there is an increase in radio emission during the epoch when the Type-C QPOs resulted in a dip in the coherence. Even the non-linear variability patterns at the QPO frequencies, explored through bicoherence studies, have been found to be correlated with spectral states and jet properties in BH LMXBs (see e.g., Arur & Maccarone, 2022, and references therein). These findings suggest that, apart from the main Comptonizing medium, the base of the jet could also act as an additional medium for Comptonization. Such hybrid geometries for the corona have been suggested by some recent spectral-timing studies (e.g., Garc´ıa et al., 2021; Méndez et al., 2022; Rout et al., 2023).